The side of a square is 10cm. How many times will the new perimeter become if the side of the square is doubled?
(A) 2 times (B) 4 times (C) 6 times (D) 8 times

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (A) is right
As we know that, perimeter of square = side × 4

= 10 × 4
= 40 cm

Given, side of square = 10 cm
The side of the square is doubled = 10 + 10

= 20 cm
Therefore, the new perimeter become 2 times the side of the square is doubled.


Thanks